# On stable assignments generated by choice functions of mixed type

### v1

@@ 2 Definitions, settings, backgrounds
$\bullet$  The stability problem SMP that we deal with is defined as follows. There is given a bipartite graph $G=(V,E)$ in which the set $V$ of vertices is partitioned into two subsets (color classes) $F$ and $W$, whose elements are conditionally called firms and workers, respectively. The edges $e\in E$ are equipped with upper bounds, or capacities, $b(e)\in{\mathbb{R}}_{+}$, and the vertices $v\in V$ with quotes $q(v)\in{\mathbb{R}}_{+}$. We will assume that all values $b(e)$ and $q(v)$ are positive; otherwise the edge/vertex can be removed without affecting the problem. One may interpret the vertices as “agents” in a market, and the edges as possible “contracts” between them whose values are bounded by the parameters $b$ and $q$. The set of edges incident to a vertex $v\in V$ is denoted by $E_{v}$.
For each vertex $v\in V$, there are preferences on the edges in $E_{v}$; they are described by a weak linear order, a transitive relation $\geq_{v}$ on pairs in $E_{v}$ in which different edges $e,e^{\prime}$ are admitted to satisfy simultaneously both comparisons $e\geq_{v}e^{\prime}$ and $e^{\prime}\geq_{v}e$. We write $e>_{v}e^{\prime}$ if $e\geq_{v}e^{\prime}$ but not $e^{\prime}\geq_{v}e$, which means that from the viewpoint of the vertex (“agent”) $v$, the edge $e$ is preferred to (or better than) the edge $e^{\prime}$; whereas if both $e\geq_{v}e^{\prime}$ and $e^{\prime}\geq_{v}e$ take place, then the edges $e,e^{\prime}$ are regarded by $v$ as equivalent, which is denoted as $e\sim_{v}e^{\prime}$.
It is useful for us to associate the weak order $\geq_{v}$ with partition $\Pi_{v}=(\pi^{1},\pi^{2},\ldots,\pi^{k})$ of the set $E_{v}$ such that $e\in\pi^{i}$ and $e^{\prime}\in\pi^{j}$ satisfy $e>_{v}e^{\prime}$ if $i<j$, and $e\sim_{v}e^{\prime}$ if $i=j$. We refer to $\Pi_{v}$ as a cortege, and to its members $\pi^{i}$ as ties, and may write $\pi^{1}>_{v}\pi^{2}>_{v}\cdots>_{v}\pi^{k}$. When $i<j$, we say that the tie $\pi^{i}$ is more preferable, or better, or earlier than $\pi^{j}$ (and $\pi^{j}$ is less preferable, or worse, or later than $\pi^{i}$).
$\bullet$  The set of admissible assignments for $(G,b)$ (without involving quotas $q$) is defined to be the box :=${\cal B}:=\{x\in{\mathbb{R}}_{+}^{E}\colon x\leq b\}$. For a vertex $v\in V$, the restriction of $x\in{\cal B}$ to the edge set $E_{v}$ is denoted by $x_{v}$; the set of these restrictions is denoted as ${\cal B}_{v}$. On this set, a choice function (CF) $C_{v}:{\cal B}_{v}\to{\cal B}_{v}$ acts as follows: for each $z\in{\cal B}_{v}$, $C_{v}(z)\leq z$, and $C_{v}$ is agreeable with the order $\geq_{v}$ and the quota $q(v)$, which means that:
@@
      if $|z|\geq q(v)$, then in the cortege $\Pi_{v}$, we take the tie $\pi^{i}$ such that $z(\pi^{1}\cup\cdots\cup\pi^{i-1})<q(v)\leq z(\pi^{1}\cup\cdots\cup\pi^{i})$, and define:
      $C_{v}(z)(e):=\begin{cases}z(e)&\mbox{if $e\in\pi^{j}$, $j<i$,}\\ 0&\mbox{if $e\in\pi^{j}$, $j>i$,}\\ r\wedge z(e)&\mbox{if $e\in\pi^{i}$;}\end{cases}$ (2.2)
@@
(Hereinafter, for real-valued functions $a,b$ on a set $S$, we use the standard notation $a\wedge b$ and $a\vee b$ for the functions taking values $\min\{a(e),b(e)\}$ and $\max\{a(e),b(e)\}$, $e\in S$, respectively. If $S^{\prime}$ is a finite subset of $S$, we write $a(S^{\prime})$ for the value $\sum(a(e):e\in S^{\prime})$. When $S$ is finite, $a(S)$ is also denoted as $|a|$.)
The tie $\pi^{i}$ figured in case ((2.1))(b) is called critical and denoted as $\pi^{\rm c}_{v}(z)$. The set of ties $\pi^{i}\in\Pi_{v}$ better (resp. worse) than $\pi^{\rm c}_{v}(z)$, i.e. $\pi^{i}>_{v}\pi^{\rm c}_{v}(z)$ (resp. $\pi^{i}<_{v}\pi^{\rm c}_{v}(z)$) is denoted by $\Pi^{\rm beg}_{v}(z)$ (resp. by $\Pi^{\rm end}_{v}(z)$). Note that $C_{v}(z)$ takes zero values on all edges of ties in $\Pi^{\rm end}_{v}(z)$.
We call $C_{v}$ defined as in ((2.1)) a choice function of the mixed type, or simply a mixed CF. Here the term “mixed” is justified by the fact that such a CF appears as a direct generalization of two cases known in the literature, pointed out as examples in [1].
Example 1. Let $\geq_{v}$ be a (strong) linear order on $E_{v}$; in other words, each tie in the cortege $\Pi_{v}$ consists of a unique element. Then the CFs $C_{v}$ as above are exactly those arising in the stable allocation problem of Bai̋ou and Balinsky [2] (denoted as SAP).
@@
$\bullet$  It is not difficult to check that a CF $C=C_{v}$ of mixed type satisfies the properties of consistence and persistence, in terminology of [1]; namely, for $z,z^{\prime}\in{\cal B}_{v}$, there hold:
  • (i)
@@
  • (ii)
    If $z\geq z^{\prime}$, then $C(z)\wedge z^{\prime}\leq C(z^{\prime})$ (persistence).
(One more axiom imposed in [1] is the size monotonicity one; its strengthening, the quota filling axiom, saying that $|z|\geq q(v)$ implies $|C(z)|=q(v)$, is automatically valid in our case.)
In particular, (i) and (ii) provide the equality $C(C(z))=C(z)$ for all $z\in{\cal B}_{v}$. A function (assignment) $z\in{\cal B}_{v}$ such that $z=C(z)$ is called rational; the set of these is denoted by ${\cal A}_{v}$. Following a general model of Alkan and Gale, the set ${\cal A}_{v}$ is endowed with the binary relation of “revealed preference” $\succeq$, where for $z,z^{\prime}\in{\cal A}_{v}$, we write $z\succeq z^{\prime}$ and say that $z$ is preferred to (or better than) $z^{\prime}$ if
$C(z\vee z^{\prime})=z.$
@@
For further needs, it is useful to specify the features of relation $\succeq$ in our case. We say that $z\in{\cal A}_{v}$ is fully filled if $|z|=q(v)$, and deficit if $|z|<q(v)$. For a fully filled $z$, in the critical tie $\pi^{\rm c}_{v}(z)$ we distinguish the subset
@@
called the head of $z$. The set of edges occurring in the ties $\pi^{1},\pi^{2},\ldots,\pi^{\rm c}_{v}(x)$, except for those in $H_{v}(z)$, is denoted by $T_{v}(z)$ and called the tail of $z$. When $z$ is deficit, we formally define $H_{v}(z):=\emptyset$ and $T_{v}(z):=E_{v}$. One can check that
  1. (2.4)
    if both $z,z^{\prime}\in{\cal A}_{v}$ are fully filled, and $\pi^{i}$ is the critical tie $\pi^{\rm c}_{v}(z)$ for $z$, then $z\succeq z^{\prime}$ if and only if $z(e)\geq z^{\prime}(e)$ holds for all edges $e$ in the tail $T_{v}(z)$ (implying that $\pi^{\rm c}_{v}(z)\geq\pi^{\rm c}_{v}(z^{\prime})$, and if $\pi^{\rm c}_{v}(z^{\prime})=\pi^{\rm c}_{v}(z)$, then $H_{v}(z^{\prime})\subseteq H_{v}(z)$ and $r^{z^{\prime}}\geq r^{z}$).
@@
Definition 1. For a vertex $v$ and a rational assignment $z\in{\cal A}_{v}$, an edge $e\in E_{v}$ is called interesting (in the sense, that an increase at $e$ could promise the existence of an assignment better than $z$) if there exists $z^{\prime}\in{\cal B}_{v}$ such that $z^{\prime}(e)>z(e)$, $z^{\prime}(e^{\prime})=z(e^{\prime})$ for all $e^{\prime}\neq e$, and $C_{v}(z^{\prime})(e)>z(e)$. (In terminology of [1], such an $e$ is called non-satiated.)
For a mixed CF $C_{v}$, one can see that:
@@
Indeed, let $z^{\prime}$ be obtained from $z$ by increasing the value at $e$ by a sufficiently small amount $\varepsilon>0$. Then $z^{\prime}\in{\cal B}_{v}$. In case $|z|<q(v)$, we have $e\in T_{v}(z)$, $z^{\prime}\in{\cal A}_{v}$, and $e$ is interesting under $z$ (since $C(z^{\prime})(e)=z^{\prime}(e)>z(e)$). Now let $|z|=q(v)$. If $e\in T_{v}(z)$, then $C_{v}$ does not change $z^{\prime}$ within the tail $T_{v}(z)$ and decreases the values within $H_{v}(z)$ by $\varepsilon/|H_{v}(z)|$ (cf. (2.2)); therefore, $C_{v}(z^{\prime})(e)=z^{\prime}(e)>z(e)$, and $e$ is interesting under $z$ again. And if $e\notin T_{v}(z)$ (i.e., $e$ is in $H_{v}(z)$ or in a tie in $\Pi^{\rm end}_{v}(z)$), then $C_{v}$ applied to $z^{\prime}$ decreases $z^{\prime}(e)$ by $\varepsilon$. So $C(z^{\prime})=z$ and $e$ is not interesting under $z$.
$\bullet$  Now we leave “individual agents” and come to assignments on the whole edge set $E$. An admissible assignment $x\in{\cal B}$ is regarded as rational if it is so for all agents, i.e., $x_{v}\in{\cal A}_{v}$ for all $v\in V$. The set of these is denoted by ${\cal A}$. So any rational assignment $x\in{\cal A}$ respects both the upper bounds $b$ and the quotas $q$.
In what follows an edge connecting vertices $f\in F$ and $w\in W$ is denoted as $fw$.
@@
We denote the set of stable assignments by ${\cal S}={\cal S}(G,b,q,\geq)$. Using ((2.5)), we can characterize blocking edges for rational assignments as follows:
@@
    for $x\in{\cal A}$, an edge $e=fw\in E$ with $x(e)<b(e)$ is blocking if and only if it belongs to both tails $T_{f}(x)$ and $T_{w}(x)$ (where $T_{v}(x)$ means $T_{v}(x_{v})$).
For stable $x,y\in{\cal S}$, when the relation $x_{f}\succeq y_{f}$ takes place for all firms $f\in F$, we write $x\succeq_{F}y$; similarly, when $x_{w}\succeq y_{w}$ holds for all workers $v\in W$, we write $x\succeq_{W}y$. According to general results on stability for bipartite graphs (subject to the size monotonicity axiom) shown by Alkan and Gale [1], the following is valid:
@@
      (polarity) the order $\succeq_{F}$ is converse to $\succeq_{W}$, namely: if $x\succeq_{F}y$ then $y\succeq_{W}x$, and vice versa;
@@
We denote the minimal and maximal elements in $({\cal S},\succeq_{W})$ by $x^{\rm min}$ and $x^{\rm max}$, respectively (so the former is the best, and the latter is the worst w.r.t. $F$).
Let us say that a vertex $v$ is fully filled if the function (restriction) $x_{v}$ is fully filled for $x\in{\cal S}$ (i.e., $|x_{v}|=q(v)$), and deficit otherwise (i.e., $|x_{v}|<q(v)$); this does not depend on the choice of stable $x$, by ((2.7))(b). We will use one more property, strengthening (b) for deficit vertices, which is shown in [1, Corollary 3] for the quota filling case (and therefore, applicable to our mixed model); it says that:

### v2

@@ 2 Definitions, settings, backgrounds
$\bullet$  The stability problem SMP that we deal with is defined as follows. There is given a bipartite graph $G=(V,E)$ in which the set $V$ of vertices is partitioned into two subsets (color classes) $F$ and $W$ whose elements are conditionally called firms and workers, respectively. The edges $e\in E$ are equipped with upper bounds, or capacities, $b(e)\in{\mathbb{R}}_{+}$, and the vertices $v\in V$ with quotes $q(v)\in{\mathbb{R}}_{+}$. We may assume that all values $b(e)$ and $q(v)$ are positive; otherwise the edge/vertex can be removed without affecting the problem. The set of edges incident to a vertex $v\in V$ is denoted by $E_{v}$.
For each vertex $v\in V$, there are preferences on the edges in $E_{v}$; they are described by a weak linear order, a transitive relation $\geq_{v}$ on pairs in $E_{v}$ in which different edges $e,e^{\prime}$ are admitted to satisfy simultaneously both comparisons $e\geq_{v}e^{\prime}$ and $e^{\prime}\geq_{v}e$. We write $e>_{v}e^{\prime}$ if $e\geq_{v}e^{\prime}$ but not $e^{\prime}\geq_{v}e$, which means that from the viewpoint of the vertex (“agent”) $v$, the edge (“contract”) $e$ is preferred to, or better than, the edge $e^{\prime}$; whereas if both $e\geq_{v}e^{\prime}$ and $e^{\prime}\geq_{v}e$ take place, then the edges $e,e^{\prime}$ are regarded by $v$ as equivalent, which is denoted as $e\sim_{v}e^{\prime}$.
Equivalently, the weak order $\geq_{v}$ can be described via the ordered partition $\Pi_{v}=(\pi^{1},\pi^{2},\ldots,\pi^{k})$ of the set $E_{v}$, where $e\in\pi^{i}$ and $e^{\prime}\in\pi^{j}$ satisfy $e>_{v}e^{\prime}$ if $i<j$, and $e\sim_{v}e^{\prime}$ if $i=j$. We refer to $\Pi_{v}$ as a cortege, and to its members $\pi^{i}$ as ties, and may write $\pi^{1}>_{v}\pi^{2}>_{v}\cdots>_{v}\pi^{k}$. When $i<j$, we say that the tie $\pi^{i}$ is more preferable, or better, or earlier than $\pi^{j}$ (and $\pi^{j}$ is less preferable, or worse, or later than $\pi^{i}$).
Next we associate the weak order $\geq_{v}$ with an appropriate choice function $C_{v}$.
$\bullet$  The set of admissible assignments for $(G,b)$ (without involving quotas $q$) is defined to be the box :=${\cal B}:=\{x\in{\mathbb{R}}_{+}^{E}\colon x\leq b\}$. For a vertex $v\in V$, the restriction of $x\in{\cal B}$ to the edge set $E_{v}$ is denoted by $x_{v}$; the set of these restrictions is denoted as ${\cal B}_{v}$. The choice function (CF) $C_{v}$ is a map of ${\cal B}_{v}$ into itself such that: for each $z\in{\cal B}_{v}$, there holds $C_{v}(z)\leq z$ (as usual for choice functions), and $C_{v}$ is agreeable with the order $\geq_{v}$ and the quota $q(v)$ in the following way:
@@
      if $|z|\geq q(v)$, then in the cortege $\Pi_{v}$, we take the tie $\pi^{i}$ such that $z(\pi^{1}\cup\cdots\cup\pi^{i-1})<q(v)\leq z(\pi^{1}\cup\cdots\cup\pi^{i})$ and define:
      $C_{v}(z)(e):=\begin{cases}z(e)&\mbox{if $e\in\pi^{j}$, $j<i$,}\\ 0&\mbox{if $e\in\pi^{j}$, $j>i$,}\\ r\wedge z(e)&\mbox{if $e\in\pi^{i}$,}\end{cases}$ (2.2)
@@
(Hereinafter, for real-valued functions $a,b$ on a set $S$, we use the standard notation $a\wedge b$ and $a\vee b$ for the functions on $S$ taking values $\min\{a(e),b(e)\}$ and $\max\{a(e),b(e)\}$, $e\in S$, respectively. If $S^{\prime}$ is a finite subset of $S$, we write $a(S^{\prime})$ for $\sum(a(e):e\in S^{\prime})$. When $S$ is finite, $\sum(|a(e)|\colon e\in S)$ is denoted as $|a|$. So $a(S)=|a|$ if $a$ is nonnegative.)
The tie $\pi^{i}$ figured in case ((2.1))(b) is called critical and denoted as $\pi^{\rm c}_{v}(z)$. The set of ties $\pi^{i}\in\Pi_{v}$ better than $\pi^{\rm c}_{v}(z)$ (i.e. $\pi^{i}>_{v}\pi^{\rm c}_{v}(z)$) is denoted by $\Pi^{\rm beg}_{v}(z)$, and the set of ties $\pi^{j}$ worse than $\pi^{\rm c}_{v}(z)$ (i.e. $\pi^{\rm c}_{v}(z)>\pi^{j}$) by $\Pi^{\rm end}_{v}(z)$). Also we denote the set of edges occurring in the ties in $\Pi^{\rm beg}_{v}(x)$ (resp. $\Pi^{\rm end}_{v}(x)$) by $E^{\rm beg}_{v}(x)$ (resp. $E^{\rm end}_{v}(x)$). Note that $C_{v}(z)$ takes zero values on all edges in $E^{\rm end}_{v}(z)$.
We call $C_{v}$ defined as in ((2.1)) a choice function of mixed type, or simply a mixed CF. Here the term “mixed” is justified by the fact that $C_{v}$ of this sort arises as a generalization of two known cases of choice functions, pointed out as examples in [1].
Example 1. Let $\geq_{v}$ be a strong linear order on $E_{v}$; in other words, each tie in the cortege $\Pi_{v}$ consists of a unique element. Then $C_{v}$ as above is a CF associated with the stable allocation problem of Bai̋ou and Balinsky [2] (denoted as SAP).
@@
$\bullet$  Like the choice functions in the above examples, a CF of mixed type satisfies the standard axioms of consistence and persistence (in terminology of [1]); let us check this for completeness of our description.
Lemma 2.1
$C=C_{v}$ defined in ((2.1)) satisfies the following conditions: for $z,z^{\prime}\in{\cal B}_{v}$,
  • (A1)
@@
  • (A2)
    if $z\geq z^{\prime}$, then $C(z)\wedge z^{\prime}\leq C(z^{\prime})$ (persistence, or substitutability).
Proof  Let us check (A1). In case $|z|\leq q(v)$, we have $C(z)=z$, and $C(z^{\prime})=C(z)$ trivially follows. And in case $|z|>q(v)$, consider the tie $\pi^{i}$ as in ((2.1))(b). Let $H:=\{e\in\pi^{i}\colon z(e)\geq r\}$. Then (2.2) shows that $C(z)(e)=z(e)$ for all $e\in E^{\rm beg}_{v}(z)\cup(\pi^{i}-H)$;  $C(z)(e)=r$ for all $e\in H$; and $C(z)(e)=0$ for all $e\in E^{\rm end}_{v}(z)$. Now $z\geq z^{\prime}\geq C(z)$ implies $z^{\prime}(e)=z(e)$ for all $e\in E^{\rm beg}_{v}(z)\cup(\pi^{i}-H)$, and $z^{\prime}(e)\geq r$ for all $e\in H$, whence $C(z^{\prime})=C(z)$ easily follows.
Next we check (A2). If $|z^{\prime}|\leq q(v)$, then $C(z^{\prime})=z^{\prime}\geq z^{\prime}\wedge C(z)$, as required. Now let $|z^{\prime}|>q(v)$. Then $z\geq z^{\prime}$ implies $|z|>q(v)$. Define $\pi^{i}$ and $r$ as in ((2.1))(b) for $z$, and define $\pi^{i^{\prime}}$ and $r^{\prime}$ as in ((2.1))(b) for $z^{\prime}$. Let $H:=\{e\in\pi^{i}\colon z(e)\geq r\}$ and $H^{\prime}:=\{e\in\pi^{i^{\prime}}\colon z^{\prime}(e)\geq r^{\prime}\}$. From $z\geq z^{\prime}$ it follows that $i\leq i^{\prime}$. Suppose that $i<i^{\prime}$. Then for $e\in E^{\rm beg}_{v}(z)\cup(\pi^{i}-H)$, we have $C(z)(e)\wedge z^{\prime}=z(e)\wedge z^{\prime}(e)=z^{\prime}(e)=C(z^{\prime})(e)$, and for $e\in H$, we have $C(z)(e)\wedge z^{\prime}(e)=r\wedge z^{\prime}(e)\leq C(z^{\prime})(e)$. It follows that $C(z)\wedge z^{\prime}\leq C(z^{\prime})$ (taking into account that $C(z)(e^{\prime})=0$ for $e^{\prime}\in E^{\rm end}_{v}(z)$).
Finally, suppose that $i=i^{\prime}$. Using $z\geq z^{\prime}$ and $|C(z)|=|C(z^{\prime})|=q(v)$, one can conclude that $r^{\prime}\geq r$ and $H^{\prime}\subseteq H$. As before, for $e\in E^{\rm beg}_{v}(z)\cup(\pi^{i}-H)$, we have $C(z)(e)\wedge z^{\prime}=z(e)\wedge z^{\prime}(e)=C(z^{\prime})(e)$. And for $e\in H$, we have $C(z)(e)\wedge z^{\prime}(e)=r\wedge z^{\prime}(e)\leq r^{\prime}\wedge z^{\prime}(e)\leq C(z^{\prime})(e)$ (using $r^{\prime}\geq r$ and $H^{\prime}\subseteq H$). As a result, we obtain the desired relation $C(z)\wedge z^{\prime}\leq C(z^{\prime})$. $\Box$
In particular, (A1) provides the equality $C(C(z))=C(z)$ for all $z\in{\cal B}_{v}$. The set of assignments $z\in{\cal B}_{v}$ such that $C(z)=z$ is denoted by ${\cal A}_{v}$ (referring to ${\cal A}_{v}$ as the set of stationary assignments on $E_{v}$). Following a general model of Alkan and Gale, the set ${\cal A}_{v}$ is endowed with the binary relation of “revealed preferences” $\succeq\,=\,\succeq_{v}$, where for $z,z^{\prime}\in{\cal A}_{v}$, one writes $z\succeq z^{\prime}$ and says that $z$ is preferred to, or better than, $z^{\prime}$ if
$C(z\vee z^{\prime})=z.$ (2.3)
@@
One more axiom imposed on CFs in Alkan–Gale’s theory is:
  • (A3)
    for $z,z^{\prime}\in{\cal B}_{v}$, if $z\geq z^{\prime}$, then $|C_{v}(z)|\geq|C_{v}(z^{\prime})|$ (size monotonicity).
A sharper form of (A3) is the quota acceptability axiom which says that $|C(z)|=\min\{|z|,q(v)\}$; this is automatically valid for the mixed CF $C_{v}$.
$\bullet$  For further needs, it is useful to specify the features of relation $\succeq$ in our case. We say that $z\in{\cal A}_{v}$ is quota filling if $|z|=q(v)$, and deficit if $|z|<q(v)$. For a quota filling $z$, in the critical tie $\pi^{\rm c}_{v}(z)=\pi^{i}$, we distinguish the subset
@@
called the head of $z$. The set of edges occurring in $E^{\rm beg}_{v}(z)\cup(\pi^{\rm c}_{v}(z)-H_{v}(z))$ is denoted by $T_{v}(z)$ and called the tail of $z$. When $z$ is deficit, we formally define $H_{v}(z):=\emptyset$ and $T_{v}(z):=E_{v}$.
It is useful to reformulate the “revealed preference” relation for assignments in ${\cal A}_{v}$ in terms of heads and tails, as follows.
Lemma 2.2
(i) For $z,z^{\prime}\in{\cal A}_{v}$, the relation $z\succeq_{v}z^{\prime}$ holds if and only if all edges $e\in T_{v}(z)$ satisfy $z(e)\geq z^{\prime}(e)$.
(ii) if $z\succ_{v}z^{\prime}$ and both $z,z^{\prime}$ are quota filling, then $\pi^{\rm c}_{v}(z)\geq\pi^{\rm c}_{v}(z^{\prime})$, and in case $\pi^{\rm c}_{v}(z)=\pi^{\rm c}_{v}(z^{\prime})$, there hold $H_{v}(z^{\prime})\subseteq H_{v}(z)$ and $r^{z^{\prime}}>r^{z}$.
Proof  The inequalities for edges in (i) easily imply (2.3) if $|z|<q(v)$ (in view of $T_{v}(z)=E_{v}$). And if $|z|=q(v)$, then these inequalities imply $C(z\vee z^{\prime})(e)=C(z)(e)=z(e)$ for all $e\in T_{v}(z)$. Also $|z|=q(v)$ implies $|C(z)|=q(v)=|C(z\vee z^{\prime})|$. These relations give $C(z\vee z^{\prime})(e^{\prime})=C(z)(e^{\prime})$ for all $e^{\prime}\in H_{v}(z)$, and (2.3) follows. The converse implication in (i) is easy as well.
To see (ii), let $\Pi_{v}=(\pi^{1}>\cdots>\pi^{k})$ and $\pi^{\rm c}_{v}(z)=\pi^{i}$. Using the inequalities $z(e)\geq z^{\prime}(e)$ for $e\in T_{v}(z)$ (valid by (i)), we have
$z^{\prime}(H_{v}(z))+z^{\prime}(E^{\rm end}_{v}(z))=q(v)-z^{\prime}(T_{v}(z))\geq q(v)-z(T_{v}(z))=z(H_{v}(z)).$
It follows that $\pi^{\rm c}_{v}(z^{\prime})\in\{\pi^{i},\ldots,\pi^{k}\}$, i.e. $\pi^{\rm c}_{v}(z^{\prime})\leq\pi^{i}=\pi^{\rm c}_{v}(z)$. If $\pi^{\rm c}_{v}(z^{\prime})=\pi^{\rm c}_{v}(z)$ takes place, then $z^{\prime}(H_{v}(z))\geq z(H_{v}(z))=r^{z}|H_{v}(z)|$ (since $z^{\prime}(E^{\rm end}_{v}(z))=0$). Then
$r^{z^{\prime}}\geq\max\{z^{\prime}(e)\colon e\in H_{v}(z)\}\geq r^{z}.$
This and $z^{\prime}(e)\leq z(e)<r^{z}$ for all $e\in\pi^{\rm c}_{v}(z)-H_{v}(z)\subseteq T_{v}(z)$ imply $H_{v}(z^{\prime})\subseteq H_{v}(z)$ and $r^{z^{\prime}}>r^{z}$ (the equality $r^{z^{\prime}}=r^{z}$ would lead to $z^{\prime}=z$, contrary to $z^{\prime}\prec_{v}z$.) $\Box$
@@
Definition 1. For a vertex $v$ and an assignment $z\in{\cal A}_{v}$, an edge $e\in E_{v}$ is called interesting (in the sense that an increase at $e$ could cause the appearance of an assignment better than $z$) if there exists $z^{\prime}\in{\cal B}_{v}$ such that $z^{\prime}(e)>z(e)$, $z^{\prime}(e^{\prime})=z(e^{\prime})$ for all $e^{\prime}\neq e$, and $C_{v}(z^{\prime})(e)>z(e)$. (In terminology of [1], such an $e$ is called non-satiated.)
For the mixed CF $C_{v}$, one can see that:
@@
Indeed, let $z^{\prime}$ be obtained from $z$ by increasing the value at $e$ by a small amount $\varepsilon>0$. Then $z^{\prime}\in{\cal B}_{v}$. In case $|z|<q(v)$, we have $e\in T_{v}(z)$, $z^{\prime}\in{\cal A}_{v}$, and $e$ is interesting under $z$ (since $C(z^{\prime})(e)=z^{\prime}(e)>z(e)$). Now let $|z|=q(v)$. If $e\in T_{v}(z)$, then $C_{v}$ does not change $z^{\prime}$ within the tail $T_{v}(z)$ and decreases the value on each edge in $H_{v}(z)$ by $\varepsilon/|H_{v}(z)|$ (cf. (2.2)); therefore, $C_{v}(z^{\prime})(e)=z^{\prime}(e)>z(e)$, and $e$ is interesting under $z$. And if $e\notin T_{v}(z)$ (i.e., $e$ is in $H_{v}(z)$ or in $E^{\rm end}_{v}(z)$), then $C_{v}$ applied to $z^{\prime}$ decreases $z^{\prime}(e)$ by $\varepsilon$. So $C(z^{\prime})=z$ and $e$ is not interesting under $z$.
$\bullet$  Now we leave “individual agents” and come to assignments on the whole edge set $E$. We define ${\cal A}$ to be the set of admissible assignments $x\in{\cal B}$ such that for all $v\in V$, the restriction $x_{v}$ of $x$ to $E_{v}$ belongs to ${\cal A}_{v}$ (referring to ${\cal A}$ as the set of fully stationary assignments on $E$). So in our case, ${\cal A}$ consists of the assignments respecting both the capacities $b$ and the quotas $q$.
In what follows, an edge connecting vertices $f\in F$ and $w\in W$ is denoted as $fw$.
@@
We denote the set of stable assignments by ${\cal S}={\cal S}(G,b,q,\geq)$. Using ((2.5)), we can characterize the blocking edges for assignments in ${\cal A}$ as follows:
@@
    for $x\in{\cal A}$, an edge $e=fw\in E$ is blocking if and only if $x(e)<b(e)$ and $e$ belongs to both tails $T_{f}(x)$ and $T_{w}(x)$ (where $T_{v}(x)$ means $T_{v}(x_{v})$).
For stable $x,y\in{\cal S}$, when the relation $x_{f}\succeq_{f}y_{f}$ takes place for all “firms” $f\in F$, we write $x\succeq_{F}y$; similarly, when $x_{w}\succeq_{w}y_{w}$ holds for all “workers” $w\in W$, we write $x\succeq_{W}y$. According to general results on stability for bipartite graphs shown by Alkan and Gale in [1], subject to axioms (A1)–(A3), the following properties are valid:
@@
      (polarity) the order $\succeq_{F}$ is converse to $\succeq_{W}$, namely: if $x\succeq_{F}y$, then $y\succeq_{W}x$, and vice versa;
@@
We denote the minimal and maximal elements in $({\cal S},\succeq_{W})$ by $x^{\rm min}$ and $x^{\rm max}$, respectively (so the former is the best, and the latter is the worst w.r.t. the part $F$).
Let us say that a vertex $v$ is fully filled by $x\in{\cal S}$ if the restriction function $x_{v}$ is quota filling (i.e. $|x_{v}|=q(v)$), and deficit otherwise (i.e. $|x_{v}|<q(v)$); this does not depend on the choice of stable $x$, by ((2.7))(b). We will use one more property shown in [1, Corollary 3] for the quota-acceptable case (and therefore, applicable to our mixed model); it strengthens ((2.7))(b) for deficit vertices, saying that